OpenStudy (anonymous):

If x^2 + 2xy+ y^2 = 25 what is the value of 4(x+y)^2?

OpenStudy (anonymous):

@Loser66 @mathsuxx123 please help

OpenStudy (anonymous):

@Zarkon please help too

OpenStudy (mertsj):

Factor the left side.

OpenStudy (mertsj):

Can you do that?

OpenStudy (anonymous):

not rlly i know it would be with x+y x+y but where would the 2 go?

OpenStudy (anonymous):

(a+b)^2=a^2+2ab+b^2

OpenStudy (mertsj):

(x+y)^2=x&2+2xy+y^2

OpenStudy (anonymous):

oh okay and then what?

OpenStudy (mertsj):

So the factors are (x+y)^2

OpenStudy (anonymous):

so 25 times 4?

OpenStudy (anonymous):

yep

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

thanks!!!

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

wait i have one more question!

OpenStudy (mertsj):

ok

OpenStudy (anonymous):

If x & y are both odd integers, which would also be an integer?? (x^2 - y^2)/(2) (x^2 - y^2)/(3) xy/2 xy/3 OR xy/6 ???

OpenStudy (mertsj):

Did you try any of the answers?

OpenStudy (anonymous):

i think its xy/3 but i dont really know how to approach this

OpenStudy (mertsj):

You could say that an odd number is one more than an even number and could be represented by 2n+1, 2n+3 and so on.

OpenStudy (anonymous):

so is xy/3 right?

OpenStudy (mertsj):

no. 5 x 7 = 35 and 35/3 is not an integer

OpenStudy (anonymous):

its x^2-y^2/3 ? :)

OpenStudy (mertsj):

(2n+1)^2-(2n+3)^2= 4n^2+4n+1-(4n^2+12n+9)= -8n-8

OpenStudy (mertsj):

Looks to me like that is divisible by 2

OpenStudy (anonymous):

OH yeah i meant 2

OpenStudy (anonymous):

thank you! :)

OpenStudy (mertsj):

yw